# FinVet: A Collaborative Framework of RAG and External Fact-Checking Agents for Financial Misinformation Detection

### v1

@@ V-C Baseline Evaluation
FinVet is evaluated against three baseline approaches: RAG (LLaMA-3.3-70B-Instruct), and Zero-shot and Chain-of-Thought (GPT-3.5-turbo). Table I shows the results. FinVet achieves an F1 score of 0.85, demonstrating substantial improvements over all baselines. The retrieval-augmented baseline, RAG (LLaMA-3.3-70B), achieves an F1 of 0.62, with notably higher precision (0.69) than recall (0.56), suggesting conservative prediction behavior when using retrieval alone. FinVet improves upon this by 37% in F1 score while achieving more balanced precision (0.86) and recall (0.84).
@@ V-D Ablation Study
  • RAG (LLaMA-3.3-70B): Single RAG pipeline using LLaMA-3.3-70B with vector store retrieval, excluding fact-checking components.
  • RAG (Mixtral-8x7B): Single RAG pipeline using Mixtral-8x7B with vector store retrieval, excluding fact-checking components.
Table II shows the ablation results. The individual RAG pipelines achieve F1 scores of 0.49 (Mixtral-8x7B) and 0.62 (LLaMA-3.3-70B), providing moderate baseline performance. The Google Fact Check API alone also achieves 0.62, though limited by its coverage of claims. Combining the API with LLM reasoning in the Fact-Check Pipeline improves performance to 0.77. Finally, FinVet’s full integration of both RAG pipelines and fact-checking achieves 0.85, which is a 10.4% relative improvement over the best individual component (Fact-Check Pipeline) and a 37% improvement over the standalone RAG approaches. These empirical results demonstrate that, for financial claim verification, the proposed hybrid architecture integrating dual RAG pipelines with fact-checking capabilities significantly outperform each constituent verification method when deployed independently.
As shown in Table I, the FinVet system outperforms the zero-shot LLM baseline by a significant margin, improving F1 score by 36%. This result underscores the limitations of parametric-only reasoning in high-stakes, fact-sensitive financial contexts and highlights the importance of architectural integration across diverse verification pathways.
TABLE I: Performance Comparison of FinVet and Baseline Approaches
@@
TABLE II: Ablation Study of FinVet Components
@@ VI Ethical Considerations
    False Positives and Negatives: No detection system is perfect. False positives could incorrectly flag legitimate financial information, potentially harming honest actors. In contrast, false negatives could allow harmful misinformation to spread. In this research, uncertainty handling has been implemented through explicit ”NEI” verdicts to mitigate unwarranted classifications, but users should be aware of these limitations. To mitigate these risks, transparency is advocated in the decision-making process, including clear documentation of pipeline biases and confidence thresholds. For future deployments, the incorporation of diverse sources of fact-checking and regular audits is recommended to ensure fairness and accountability. Interdisciplinary collaboration with ethicists and financial regulators is also encouraged to align the system with societal values and legal standards.
VII Conclusion and Future Work
This research introduced FinVet, a novel framework for financial misinformation detection that integrates RAG pipelines with external fact-checking through a confidence-weighted voting mechanism. Experimental results demonstrate that the framework achieves an F1 score of 85%, which is an 8 % improvement over the best individual pipeline. This substantial performance gain validates three core design principles: the complementary value of diverse verification pathways, the importance of external verification source integration, and the effectiveness of confidence-based verdict integration. Moreover, FinVet’s ability to generate evidence-based justifications, confidence scores, and transparent source attributions addresses key requirements for financial institutions, regulators, and oversight bodies that rely on auditable verification and decision traceability.
Future research will incorporate additional financial knowledge sources—such as regulatory filings, earnings reports, market analyses, and financial news—to further enhance domain-specific retrieval and coverage. In addition, semantic-aware evaluation metrics will be explored to assess the quality, credibility, and contextual relevance of the evidence supporting each verdict.
To further improve the accuracy and adaptability of the framework, the LLMs will be fine-tuned using financial-domain instruction datasets, user feedback loops will be implemented to continuously refine retrieval and verification quality, and multilingual support will be extended to facilitate the detection of cross-border financial misinformation more effectively.

### v2

@@ V-C Baseline Evaluation
FinVet is evaluated against three baseline approaches: RAG (LLaMA-3.3-70B-Instruct), and Zero-shot and Chain-of-Thought (GPT-3.5-turbo). Table II shows the results. FinVet achieves an F1 score of 0.85, demonstrating substantial improvements over all baselines. The retrieval-augmented baseline, RAG (LLaMA-3.3-70B), achieves an F1 of 0.62, with notably higher precision (0.69) than recall (0.56), suggesting conservative prediction behavior when using retrieval alone. FinVet improves upon this by 37% in F1 score while achieving more balanced precision (0.86) and recall (0.84).
@@ V-D Ablation Study
  • RAG (LLaMA-3.3-70B): Single RAG pipeline using LLaMA-3.3-70B, excluding fact-checking components.
  • RAG (Mixtral-8x7B): Single RAG pipeline using Mixtral-8x7B, excluding fact-checking components.
Table III shows the ablation results. The individual RAG pipelines achieve F1 scores of 0.49 (Mixtral-8x7B) and 0.62 (LLaMA-3.3-70B), providing moderate baseline performance. The Google Fact Check API alone also achieves an F1 score of 0.62. Combining the external fact checking API with LLM reasoning in the Fact-Check Pipeline improves performance to 0.77. Finally, FinVet’s full integration of both RAG pipelines and fact-checking achieves an F1 score of 0.85, which is a 10.4% relative improvement over the best individual component (Fact-Check Pipeline) and a 37% improvement over the standalone RAG approaches. These empirical results demonstrate that, for financial claim verification, the proposed hybrid architecture integrating dual RAG pipelines with fact-checking capabilities significantly outperform each constituent verification method when deployed independently.
TABLE II: Performance Comparison of FinVet and Baseline Approaches
@@
TABLE III: Ablation Study of FinVet Components
@@ VI Ethical Considerations
    False Positives and Negatives: No detection system is perfect. False positives could incorrectly flag legitimate financial information, potentially harming honest actors. In contrast, false negatives could allow harmful misinformation to spread. In this research, uncertainty handling has been implemented through explicit "NEI" verdicts to mitigate unwarranted classifications, but users should be aware of these limitations. To mitigate these risks, transparency is advocated in the decision-making process, including clear documentation of pipeline biases and confidence thresholds. For future deployments, the incorporation of diverse sources of fact-checking and regular audits is recommended to ensure fairness and accountability. Interdisciplinary collaboration with ethicists and financial regulators is also encouraged to align the system with societal values and legal standards.
VII Conclusion
This research introduced FinVet, a novel framework for financial misinformation detection that integrates RAG pipelines with external fact-checking through a confidence-weighted voting mechanism. Experimental results demonstrate that the framework achieves an F1 score of 0.85, which is a 10.4% improvement over the best individual pipeline. This substantial performance gain validates three core design principles: the complementary value of diverse verification pathways, the importance of external verification source integration, and the effectiveness of confidence-based verdict integration. Moreover, FinVet’s ability to generate evidence-based justifications, confidence scores, and transparent source attributions addresses key requirements for financial institutions, regulators, and oversight bodies that rely on auditable verification and decision traceability.
VIII Future Work
Future research will extend FinVet’s evaluation to additional financial datasets and domains—such as regulatory filings, earnings reports, and market news—to assess its generalizability across various real-world financial domains. A comprehensive analysis of computational efficiency and latency will also be conducted, including per-stage runtime and scalability under high-throughput settings. To enhance robustness, multi-fold cross-validation will be implemented with variance and confidence interval reporting to demonstrate consistency across data splits. Furthermore, statistical significance testing between the full model and its ablated variants will be performed to validate the contribution of individual components.
To further enhance the framework, future work will incorporate semantic-aware evaluation metrics, fine-tune large language models on financial-domain instruction datasets, and integrate user feedback loops to continuously refine retrieval and verification quality. Finally, reproducibility will be strengthened by expanding hyperparameter documentation and releasing controlled evaluation scripts to facilitate independent verification of results.
@@ Acknowledgment
Appendix A Prompt Design Template
This appendix presents the structured prompt template used to guide the LLMs used in FinVet during the verification process. The prompt enforces role-based reasoning, explicit evidence citation, and consistency checks to minimize unsupported generations and improve transparency.
1
2prompt = f"""You are a collaborative team of expert fact-checkers evaluating claims that may impact financial markets, consisting of five specialized roles:
3
41. Financial Markets Analyst:
5 - Analyzes economic indicators, corporate performance, and market dynamics
6 - Evaluates magnitude and plausibility of claimed financial impacts
7 - Verifies historical precedents for similar market events
8 - Assesses market psychology and reaction patterns
9
102. Political Economy Specialist:
11 - Assesses government actions, regulatory changes, and geopolitical events
12 - Validates policy transmission mechanisms to economic outcomes
13 - Evaluates legislative impacts on financial sectors
14 - Understands central bank decisions and their market effects
15
163. Quantitative Verification Expert:
17 - Fact-checks specific numbers: prices, percentages, dates, amounts
18 - Verifies corporate financial data and earnings reports
19 - Confirms market timing and causation claims
20 - Performs temporal and numerical consistency checks
21
224. Misinformation Pattern Analyst:
23 - Identifies market manipulation and coordinated disinformation campaigns
24 - Detects synthetic events, deepfakes, and false announcements
25 - Recognizes financial fraud patterns and pump-and-dump schemes
26 - Evaluates information cascade and viral spread patterns
27
285. Investigative Journalist:
29 - Verifies primary sources and attribution chains
30 - Uncovers hidden connections and conflicts of interest
31 - Applies professional fact-checking standards
32 - Traces claim origins and propagation paths
33
34Carefully analyze the following claim for accuracy and verifiability.
35CITE ONLY information from the input claim and/or Retrieved Evidence.
36Do NOT introduce external facts or unsupported information.
37
38
39Decision Guidelines:
40 - If retrieved evidence directly contradicts or confirms the claim then Use it
41 - If no retrieval but claim is verifiable through logic/known facts then Provide verdict with "Parametric Knowledge" source
42 - If neither retrieval nor parametric knowledge can verify then Return NEI
43
44
45{f"Retrieved Evidence: {metadata.get('evidence_sentence', '')}" if metadata and metadata.get('evidence_sentence') else "No retrieved evidence available"}
46{f"Retrieved Label: {metadata.get('label', '')}" if metadata and metadata.get('label') else ""}
47{f"Source: {metadata.get('hrefs', '')}" if metadata and metadata.get('hrefs') else ""}
48
49Claim: {query}
50
51Analysis Framework:
52
531. Core Claims Identification:
54 - Break down the main factual assertions
55 - Identify verifiable components (events, numbers, causation)
56 - Note important context, timeframes, and qualifiers
57
582. Step-by-Step Analysis:
59 - Systematically evaluate each claim component
60 - Cross-reference with retrieved evidence
61 - Apply relevant expertise from each team member
62 - Verify both triggering events AND claimed impacts
63
643. Evidence Assessment:
65 - Evaluate reliability of available information
66 - Identify gaps requiring NEI designation
67 - Consider potential biases or misleading elements
68 - Distinguish correlation from causation
69
704. Final Determination:
71 Based on collective analysis, label the claim as:
72 - True: Claim is accurate and verified
73 - False: Claim contains verifiable inaccuracies
74 - NEI: Insufficient information to determine veracity
75
76Output Requirements:
77- Evidence: Detailed justification based solely on available information
78- Source: {f"metadata.get('hrefs', 'Parametric Knowledge')" if metadata else "'Parametric Knowledge'"}
79- Confidence: 0.0-1.0 (0.0=complete uncertainty, 1.0=absolute certainty)
80
81Return response in EXACTLY this format:
82Label: [true/false/nei]
83Evidence: [Your detailed reasoning]
84Source: [Your source of information]
85Confidence: [0.0-1.0]"""
Listing 1: FinVet LLM Prompt